Exact Reconstruction of Damaged Color Images Using a Total

Variation Model I. Fonseca Department of Mathematical Sciences Carnegie Mellon University Pittsburgh, PA 15235 U.S.A. fonseca@andrew.cmu.edu

G. Leoni Department of Mathematical Sciences Carnegie Mellon University Pittsburgh, PA 15235 U.S.A. giovanni@andrew.cmu.edu

F. Maggi Dipartimento di Matematica “U. Dini” Università di Firenze, Firenze, Italy maggi@math.unifi.it

M. Morini SISSA, Trieste, Italy morini@sissa.it

Abstract: In this paper the reconstruction of damaged piecewice constant color images is studied using a RGB total variation based model for colorization/inpainting. In particular, it is shown that when color is known in a uniformly distributed region, then reconstruction is possible with maximal fidelity.